OpenStudy (123456789):

Prove the identity sin ^2cos^2= 1/8(1-cos 4x)

OpenStudy (anonymous):

i think ur problem is incomplete.

OpenStudy (123456789):

why?

OpenStudy (123456789):

the original was : rewrite sin ^2cos^2 using the power reducing formulas and 1/8(1-cos 4x) was the answer in the book; i dont know how to get there.

OpenStudy (anonymous):

sin^2cos^2 has no variable x and the answer includes x? how come?

OpenStudy (123456789):

sin^2xcos^2x sorry

OpenStudy (anonymous):

okay.... cos 2x = cos^2 x - sin^2 x cos 2(2x) = cos^2 (2x) - sin^2 (2x) cos 4x = cos^2 (2x) - sin^2 (2x) eqn 1 sin^2 2x + cos^2 2x = 1 cos^2 2x = 1 - sin^2 2x eeqn 2 substitute eqn 2 to eqn 1: cos 4x = (1 - sin^2 2x) - sin^2 (2x) cos 4x = 1 - 2sin^2 2x 2sin^2 2x = 1 - cos 4x sin^2 2x = (1 - cos 4x)/2 eqn 3 sin 2x = 2sin x cos x sin x cos x = (sin 2x)/2 sin^2 x cos ^2 x = (sin^2 2x)/4 eqn 4 subst eqn 3 to eqn 4 sin^2 x cos ^2 x = ((1 - cos 4x)/2)/4 sin^2 x cos ^2 x = ((1 - cos 4x)/8 proven!

OpenStudy (123456789):

Thanks!